what is the given the equation x+y=6 and the assigned values for x, find the value of y.


1
Expert's answer
2021-09-27T15:06:24-0400

Solution:

Given equation: "x+y=6"

On comparing it with general form "ax+by+c=0", we get,

"a=1,b=1,c=-6"

Assume the assigned values of "x" are 0,1,2. Now we find respective values of "y" .

Put "x=0" in "x+y=6"

"\Rightarrow0+y=6
\\ \Rightarrow y=6"

Put "x=1" in "x+y=6"

"\Rightarrow1+y=6
\\ \Rightarrow y=5"

Put "x=2" in "x+y=6"

"\Rightarrow2+y=6
\\ \Rightarrow y=4"